Continuous Interpolations of Solution Sets of Quantum Stochastic Differential Inclusions

Ezekiel Ayoola Chalmers University of Technology, Sweden

Given any finite set of trajectories of a Lipschitzian quantum stochastic differential inclusion, we show that there exists a continuous selection from the complex valued multifunction associated with the solution set of the inclusion,interpolating the matrix elements of the given trajectories